Show that the conicoid 2π‘₯2 + 2𝑦2 + π‘₯𝑦 βˆ’ 𝑦𝑧 + 𝑧π‘₯ + 2π‘₯ βˆ’ 𝑦 + 5𝑧 + 1 = 0 is

central. Hence find its centre.


1
Expert's answer
2022-02-11T12:56:25-0500

1] A conicoid, given by the equation

"ax^2+by^2+cz^2+2fyz+2gxz+2hxy+2ux+2vy+2wz+d=0"

has the point "P(x_0,\ y_0,\ z_0)" as a center if and only if;

"ax_0+hy_0+gz_0+u=0\\
hx_0+by_0+fz_0+v=0\\
gx_0+fy_0+cz_0+w=0"


Now, by comparing

"ax^2+by^2+cz^2+2fyz+2gxz+2hxy+2ux+2vy+2wz+d=0"

with the given equation

"2\ud835\udc65^2 + 2\ud835\udc66^2 + \ud835\udc65\ud835\udc66 \u2212 \ud835\udc66\ud835\udc67 + \ud835\udc67\ud835\udc65 + 2\ud835\udc65 \u2212 \ud835\udc66 + 5\ud835\udc67 + 1 = 0"

we have;

"\displaystyle
a=2,\ b=2,\ c=0,\ f=-\frac{1}{2},\ g=\frac{1}{2},\ h=\frac{1}{2},\ u=1,\ v=-\frac{1}{2},\ w=\frac{5}{2},\ d=1"

Thus,

"ax_0+hy_0+gz_0+u=0\\
hx_0+by_0+fz_0+v=0\\
gx_0+fy_0+cz_0+w=0"

"\Rightarrow"

"\displaystyle
4x_0+y_0+z_0=-2\\
x_0+4y_0-z_0=1\\
x_0-y_0+0z_0=-5"


"\displaystyle
\Rightarrow x_0=-\frac{13}{5},\ y_0=\frac{12}{5},\ z_0=6"on solving the system of equations.

Thus, the center of the conicoid is "\displaystyle
P\left(-\frac{13}{5},\ \frac{12}{5},\ 6\right)".


2] A conicoid is called a central conicoid if it has a unique center.

To check if the given conicoid has a unique center, we need to show that the system of linear equations;

"\displaystyle
4x_0+y_0+z_0=-2\\
x_0+4y_0-z_0=1\\
x_0-y_0+0z_0=-5\\"


"\Rightarrow
\begin{pmatrix}
 4 & 1 & 1\\
 1 & 4 & -1\\
 1 & -1 & 0
\end{pmatrix}

\begin{pmatrix}
 x_0 \\
 y_0 \\
 z_0
\end{pmatrix}=
\begin{pmatrix}
 -2\\
 1\\
 -5
\end{pmatrix}" , in matrix form has a unique solution (exactly one solution) by showing that the determinant of the coefficient matrix is not equal to zero. That is we need to show that

"\begin{vmatrix}
 4 & 1 & 1\\
 1 & 4 & -1\\
 1 & -1 & 0
\end{vmatrix}\neq 0"

Now,

"\begin{vmatrix}
 4 & 1 & 1\\
 1 & 4 & -1\\
 1 & -1 & 0
\end{vmatrix}=-10\neq0"

Thus the coefficient matrix is invertible and "x_0,\ y_0,\ z_0" are unique from the properties of matrix, with

"\displaystyle
(x_0,\ y_0,\ z_0)=\left(-\frac{13}{5},\ \frac{12}{5},\ 6\right)"

showing that the conicoid center is unique.